### v1

@@ 5 Numerical experiments
[…]
@@ 5.1 Comparison with SNIPAL and ciPALM
[…]
We use images from the ClassicImages class in the DOTmark collection [41], which serves as a benchmark dataset for the OT problem and its variants, to generate the QROT instance. Note that the images in the ClassicImages class consist of ten different images, each with different resolutions of $32\times 32$, $64\times 64$, $128\times 128$ and $512\times 512$. Thus, for each resolution, we can pair any two different images and compute the QROT problem, resulting in 45 QROT problems. Moreover, the cost matrix $C$ is obtained by calculating the squared Euclidean distances between pixels.
[…]
Table 1: Comparisons between ripALM, SNIPAL and ciPALM under different choices of tolerance parameters, where $\lambda=1$ and the instances are generated using images with the resolution of $64\times 64$ in the ClassicImages class from the DOTmark collection. In the table, “$\Delta_{\mathrm{res}}$” denotes the terminating $\Delta^{k}_{\mathrm{res}}$; “#” denotes the number of iterations (the total number of the Ssn iterations is given in the bracket); “time” denotes the computational time in seconds.
@@
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}$ time $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}$ time
(1, 1.1, 1.1) 7.85e-07 17 (52) 9.56 (0.01, 1.1, 1.1) 7.85e-07 17 (61) 11.12
(1, 1.1, 2.1) 7.85e-07 17 (55) 9.97 (0.01, 1.1, 2.1) 7.85e-07 17 (63) 11.36
(1, 1.1, 3.1) 7.85e-07 17 (58) 10.33 (0.01, 1.1, 3.1) 7.85e-07 17 (64) 11.57
(1, 2.1, 1.1) 7.85e-07 17 (52) 9.55 (0.01, 2.1, 1.1) 7.85e-07 17 (61) 11.12
(1, 2.1, 2.1) 7.85e-07 17 (55) 9.98 (0.01, 2.1, 2.1) 7.85e-07 17 (63) 11.36
(1, 2.1, 3.1) 7.85e-07 17 (58) 10.33 (0.01, 2.1, 3.1) 7.85e-07 17 (64) 11.56
(1, 3.1, 1.1) 7.85e-07 17 (52) 9.55 (0.01, 3.1, 1.1) 7.85e-07 17 (61) 11.13
(1, 3.1, 2.1) 7.85e-07 17 (55) 9.99 (0.01, 3.1, 2.1) 7.85e-07 17 (63) 11.36
(1, 3.1, 3.1) 7.85e-07 17 (58) 10.34 (0.01, 3.1, 3.1) 7.85e-07 17 (64) 11.57
@@
$\rho$ $\Delta_{\mathrm{res}}$ time $\rho$ $\Delta_{\mathrm{res}}$ time
0.999 7.85e-07 17 (47) 8.86 0.999 7.85e-07 17 (68) 15.49
0.99 7.85e-07 17 (47) 8.86 0.99 7.85e-07 17 (68) 15.53
0.9 7.85e-07 17 (47) 8.86 0.9 7.85e-07 17 (68) 15.17
0.7 7.85e-07 17 (47) 8.99 0.7 7.85e-07 17 (66) 14.08
0.5 7.85e-07 17 (48) 9.07 0.5 7.85e-07 17 (63) 12.95
0.3 7.85e-07 17 (49) 9.21 0.3 7.85e-07 17 (60) 11.51
0.1 7.85e-07 17 (50) 9.45 0.1 7.85e-07 17 (57) 10.65
0.05 7.85e-07 17 (51) 9.61 0.05 7.85e-07 17 (56) 10.51
0.01 7.85e-07 17 (53) 9.94 0.01 7.85e-07 17 (58) 10.76
@@
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}$ time $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}$ time
(1, 1.1, 1.1) 6.43e-07 19 (92) 28.48 (0.01, 1.1, 1.1) 6.43e-07 19 (101) 30.04
(1, 1.1, 2.1) 6.43e-07 19 (95) 28.91 (0.01, 1.1, 2.1) 6.43e-07 19 (103) 30.31
(1, 1.1, 3.1) 6.43e-07 19 (98) 29.26 (0.01, 1.1, 3.1) 6.43e-07 19 (105) 30.50
(1, 2.1, 1.1) 6.43e-07 19 (92) 28.49 (0.01, 2.1, 1.1) 6.43e-07 19 (101) 30.06
(1, 2.1, 2.1) 6.43e-07 19 (95) 28.91 (0.01, 2.1, 2.1) 6.43e-07 19 (103) 30.28
(1, 2.1, 3.1) 6.43e-07 19 (98) 29.27 (0.01, 2.1, 3.1) 6.43e-07 19 (105) 30.47
(1, 3.1, 1.1) 6.43e-07 19 (92) 28.49 (0.01, 3.1, 1.1) 6.43e-07 19 (101) 30.02
(1, 3.1, 2.1) 6.43e-07 19 (95) 28.92 (0.01, 3.1, 2.1) 6.43e-07 19 (103) 30.29
(1, 3.1, 3.1) 6.43e-07 19 (98) 29.28 (0.01, 3.1, 3.1) 6.43e-07 19 (105) 30.48
@@
$\rho$ $\Delta_{\mathrm{res}}$ time $\rho$ $\Delta_{\mathrm{res}}$ time
0.999 6.43e-07 19 (86) 27.77 0.999 6.43e-07 19 (108) 34.70
0.99 6.43e-07 19 (86) 27.78 0.99 6.43e-07 19 (108) 34.71
0.9 6.43e-07 19 (86) 27.78 0.9 6.43e-07 19 (107) 34.11
0.7 6.43e-07 19 (87) 27.88 0.7 6.43e-07 19 (106) 33.29
0.5 6.43e-07 19 (87) 27.95 0.5 6.43e-07 19 (103) 32.10
0.3 6.43e-07 19 (88) 28.10 0.3 6.43e-07 19 (100) 30.60
0.1 6.43e-07 19 (90) 28.32 0.1 6.43e-07 19 (97) 29.66
0.05 6.43e-07 19 (91) 28.47 0.05 6.43e-07 19 (96) 29.45
0.01 6.43e-07 19 (92) 28.83 0.01 6.43e-07 19 (98) 29.73
5.2 Comparison with Gurobi and dADMM
@@
In this part of experiments, we compare our ripALM with Gurobi (version 10.0.1) and dADMM for solving large-scale QROT problems. For our ripALM, we set $\rho=0.99$ based on the numerical observations from the previous section. For Gurobi, we use its default termination conditions and set the corresponding termination tolerances as $10^{-6}$, aligning it with our tolerance. For dADMM, we initialize the penalty parameter as $\sigma_{0}=0.01\|C\|_{F}^{-1}$, and dynamically adjust it based on the primal-dual residuals, following the approach described in [50, Section 5.1], to achieve superior empirical performance. Moreover, we will terminate dADMM when it returns a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ satisfying $\Delta^{k}_{\mathrm{res}}<10^{-6}$, or its number of iterations reaches 10000.
@@
We follow [52, Section 5.1] to generate a random QROT instance. Specifically, we first generate two discrete probability distributions $\big{\{}(a_{i},\,\bm{p}_{i})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:i=1,\cdots,m\big{\}}$ and $\big{\{}(b_{j},\,\bm{q}_{j})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:j=1,\cdots,n\big{\}}$. Here, $\bm{a}:=(a_{1},\cdots\!,a_{m})^{\top}$ and $\bm{b}:=(b_{1},\cdots\!,b_{n})^{\top}$ are probabilities/weights, which are generated from the uniform distribution on the open interval (0, 1)01(0,\,1)( 0 , 1 ) and further normalized such that $\sum^{m}_{i}a_{i}=\sum^{n}_{j}b_{j}=1$. Moreover, $\{\bm{p}_{i}\}$ and $\{\bm{q}_{j}\}$ are support points whose entries are drawn from a five-component multivariate Gaussian mixture distribution, with a mean vector $(-20,10,0,10,20)^{\top}$ and a variance vector $(5,5,5,5,5)^{\top}$, using randomly generated weights. Then, the cost matrix $C$ is generated by $c_{ij}=\|\bm{p}_{i}-\bm{q}_{j}\|^{2}$ for $1\leq i\leq m$ and $1\leq j\leq n$ and normalized by dividing (element-wise) by its maximal entry.
[…]
@@
Table 3: Numerical results of Gurobi, dADMM, and ripALM. In the table, “$\Delta_{\mathrm{res}}$” denotes the terminating $\Delta^{k}_{\mathrm{res}}$; “#” denotes the number of iterations (the total number of the Ssn iterations is given in the bracket); “time” denotes the computational time.
Gurobi dADMM ripALM
$m=n$ $\Delta_{\mathrm{res}}$ time $\Delta_{\mathrm{res}}$ time $\Delta_{\mathrm{res}}$ time
@@
1000 2.55e-07 33 4.63 3.17e-05 10000 15.31 7.50e-07 18 (83) 1.65
2000 5.55e-07 38 22.26 1.39e-04 10000 100.22 5.77e-07 19 (107) 8.76
3000 1.75e-06 41 57.58 2.78e-04 10000 238.54 5.29e-07 19 (119) 22.82
4000 6.43e-07 43 118.20 5.01e-04 10000 430.89 6.85e-07 19 (129) 44.08
5000 4.71e-07 45 191.02 7.26e-04 10000 678.03 5.49e-07 20 (141) 78.06
6000 5.57e-07 47 298.67 1.08e-03 10000 977.75 5.92e-07 20 (148) 125.25
7000 6.98e-07 49 437.49 1.30e-03 10000 1332.25 6.21e-07 20 (149) 173.81
8000 7.95e-07 49 629.63 1.62e-03 10000 1740.13 7.59e-07 20 (157) 240.14
9000 5.87e-07 51 794.23 2.15e-03 10000 2202.89 6.03e-07 21 (168) 346.25
10000 6.02e-07 52 1073.12 2.43e-03 10000 2718.42 4.96e-07 21 (176) 462.52
[…]

### v2

@@ 5 Numerical experiments
Hyperparameters. Our ripALM as well as Snipal and ciPALM require appropriate choices of $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ to achieve superior performance. In our experiments, for all these algorithms, we simply set $\sigma_{k}=\min\big\{10^{4},\,\max\big\{10^{-4},\,1.5^{k}\big\}\big\}$ and $\tau_{k}\equiv 5$ for all $k\geq 0$, without employing any sophisticated parameter tuning. These choices of $\{\tau_{k}\}$ and $\{\sigma_{k}\}$ can satisfy the conditions required in Theorems 3.1 and 3.2; see also Remark 3.1. Moreover, we would like to emphasize that adopting more refined updating rules for $\sigma_{k}$ and $\tau_{k}$ is certainly possible and may further improve the numerical performance. In addition, for the Ssn method in Algorithm 2, we set $\eta=10^{-4}$, $\delta=0.5$, $\bar{\mu}=10^{-3}$ and $\mu=0.2$.
@@ 5.1 Comparison with SNIPAL and ciPALM
In this part of experiments, we compare our ripALM with Snipal [24] and ciPALM [58] for solving the QROT problem (4.1) to illustrate how different error criteria with different tolerance parameters influence the numerical performance of the inexact pALM. Note that both ripALM and ciPALM are of relative-type and only require a single tolerance parameter $\rho$, as shown in (2.5) and (2.8), respectively, while Snipal is of absolute-type and requires two summable tolerance parameter sequences $\{\varepsilon_{k}\}$ and $\{\delta_{k}\}$, as shown in (2.7). For simplicity, in our comparisons, we set $\varepsilon_{k}=\varepsilon_{0}/(k+1)^{p},\;\delta_{k}=\delta_{0}/(k+1)^{q}$ with $\varepsilon_{0}=\delta_{0}\in\big\{0.01,1\big\}$ and $p,q\in\big\{1.1,\,2.1,\,3.1\big\}$ (hence, there are 18 combinations in total) for Snipal. For ripALM and ciPALM, we consider $\rho\in\left\{0.01,0.05,0.1,0.3,0.5,0.7,0.9,0.99,0.999\right\}$ (hence, there are 9 choices).
We use images from the ClassicImages and MicroscopyImages classes in the DOTmark collection [44], which serves as a benchmark dataset for the OT problem and its variants, to generate QROT instances. Note that the images in the ClassicImages (or MicroscopyImages) class consist of 10 different images, each with different resolutions of $32\times 32$, $64\times 64$, $128\times 128$ and $512\times 512$. Thus, for each resolution, we can pair any two different images within a class and compute the QROT problem, resulting in 90 QROT problems (45 from ClassicImages and 45 from MicroscopyImages). Moreover, the cost matrix $C$ is obtained by calculating the squared Euclidean distances between pixels.
Tables 2 and 3 present the average results (over 90 instances) for $64\times 64$ resolution with the regularization parameter $\lambda\in\{1,\,0.1\}$. From the results, we see that the performance of all algorithms depends on the choices of tolerance parameters, and with proper tuning, their performances can be comparable. This is indeed expected because all the algorithms essentially employ the same “pALM + Ssn” framework and differ only in their error criteria for solving subproblems. Notably, since ripALM and ciPALM involve only a single tolerance parameter $\rho\in[0,1)$, they would be more user-friendly and easier to tune through a simple one-dimensional grid search. This supports the main motivation for developing a relative-type error criterion. Moreover, one can see that ripALM always outperforms ciPALM in terms of the total number of Ssn iterations and exhibits a more robust performance, particularly when a larger $\rho$ is chosen, leading to a noticeable reduction in computational time. This difference may be attributed to the extra correction step required by ciPALM to guarantee convergence, which tends to considerably increase the number of Ssn iterations [^4] One possible reason is that the variable obtained after the correction step may no longer serve as a suitable initial point for the Ssn method in the next iteration. needed to satisfy its error criterion (2.8), even though these extra iterations do not contribute to the progress of the pALM iterations. These observations further support the advantage of the vanilla inexact pALM framework advocated in this paper. As shown in Tables 2 and 3, our ripALM clearly exhibits greater robustness and efficiency. In addition, its performance remains relatively stable across different values of $\rho$, with slight advantage for $\rho\geq 0.5$.
Table 2: Comparisons between ripALM, Snipal and ciPALM under different choices of tolerance parameters for solving the QROT problem (4.1), where $\lambda=1$ and the instances are generated using images with the resolution of $64\times 64$ in the ClassicImages and MicroscopyImages classes from the DOTmark collection. In the table, $\Delta_{\mathrm{res}}^{\rm qrot}$” denotes the terminating $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$; “#” denotes the number of iterations (the total number of Ssn iterations is given in the bracket); “time” denotes the computational time in seconds.
@@
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
(1, 1.1, 1.1) 7.48e-07 17 (51) 8.90 (0.01, 1.1, 1.1) 7.48e-07 17 (67) 11.01
(1, 1.1, 2.1) 7.48e-07 17 (58) 9.76 (0.01, 1.1, 2.1) 7.48e-07 17 (70) 11.42
(1, 1.1, 3.1) 7.48e-07 17 (62) 10.23 (0.01, 1.1, 3.1) 7.48e-07 17 (72) 11.65
(1, 2.1, 1.1) 7.48e-07 17 (51) 8.89 (0.01, 2.1, 1.1) 7.48e-07 17 (67) 11.01
(1, 2.1, 2.1) 7.48e-07 17 (58) 9.74 (0.01, 2.1, 2.1) 7.48e-07 17 (70) 11.43
(1, 2.1, 3.1) 7.48e-07 17 (62) 10.21 (0.01, 2.1, 3.1) 7.48e-07 17 (72) 11.66
(1, 3.1, 1.1) 7.48e-07 17 (51) 8.89 (0.01, 3.1, 1.1) 7.48e-07 17 (67) 11.01
(1, 3.1, 2.1) 7.48e-07 17 (58) 9.75 (0.01, 3.1, 2.1) 7.48e-07 17 (70) 11.41
(1, 3.1, 3.1) 7.48e-07 17 (62) 10.22 (0.01, 3.1, 3.1) 7.48e-07 17 (72) 11.65
@@
$\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
0.999 7.48e-07 17 (42) 8.01 0.999 7.48e-07 17 (164) 71.95
0.99 7.48e-07 17 (42) 8.02 0.99 7.55e-07 17 (163) 71.53
0.9 7.48e-07 17 (42) 8.03 0.9 7.40e-07 17 (156) 67.33
0.7 7.48e-07 17 (42) 8.10 0.7 7.48e-07 17 (142) 58.87
0.5 7.48e-07 17 (43) 8.18 0.5 7.48e-07 17 (125) 48.21
0.3 7.48e-07 17 (44) 8.32 0.3 7.48e-07 17 (105) 36.16
0.1 7.48e-07 17 (47) 8.68 0.1 7.48e-07 17 (79) 20.00
0.05 7.48e-07 17 (49) 8.93 0.05 7.48e-07 17 (70) 14.73
0.01 7.48e-07 17 (55) 9.59 0.01 7.48e-07 17 (65) 11.30
@@
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
(1, 1.1, 1.1) 6.19e-07 20 (92) 27.72 (0.01, 1.1, 1.1) 6.19e-07 20 (110) 29.93
(1, 1.1, 2.1) 6.19e-07 20 (100) 28.60 (0.01, 1.1, 2.1) 6.19e-07 20 (113) 30.32
(1, 1.1, 3.1) 6.19e-07 20 (104) 29.11 (0.01, 1.1, 3.1) 6.19e-07 20 (115) 30.56
(1, 2.1, 1.1) 6.19e-07 20 (92) 27.74 (0.01, 2.1, 1.1) 6.19e-07 20 (110) 29.93
(1, 2.1, 2.1) 6.19e-07 20 (100) 28.63 (0.01, 2.1, 2.1) 6.19e-07 20 (113) 30.32
(1, 2.1, 3.1) 6.19e-07 20 (104) 29.10 (0.01, 2.1, 3.1) 6.19e-07 20 (115) 30.56
(1, 3.1, 1.1) 6.19e-07 20 (92) 27.72 (0.01, 3.1, 1.1) 6.19e-07 20 (110) 29.93
(1, 3.1, 2.1) 6.19e-07 20 (100) 28.60 (0.01, 3.1, 2.1) 6.19e-07 20 (113) 30.32
(1, 3.1, 3.1) 6.19e-07 20 (104) 29.09 (0.01, 3.1, 3.1) 6.19e-07 20 (115) 30.57
@@
$\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
0.999 6.20e-07 20 (81) 26.68 0.999 6.19e-07 20 (230) 111.70
0.99 6.20e-07 20 (81) 26.68 0.99 6.19e-07 20 (232) 113.24
0.9 6.20e-07 20 (82) 26.72 0.9 6.19e-07 20 (222) 105.97
0.7 6.20e-07 20 (82) 26.78 0.7 6.19e-07 20 (200) 91.75
0.5 6.20e-07 20 (83) 26.85 0.5 6.19e-07 20 (181) 79.50
0.3 6.19e-07 20 (85) 27.00 0.3 6.19e-07 20 (156) 63.09
0.1 6.19e-07 20 (88) 27.39 0.1 6.19e-07 20 (122) 41.27
0.05 6.19e-07 20 (90) 27.68 0.05 6.19e-07 20 (115) 36.27
0.01 6.19e-07 20 (96) 28.39 0.01 6.19e-07 20 (108) 31.14
5.2 Comparison with Gurobi, dADMM and ES-iALM
@@
In this part of experiments, we compare our ripALM with Gurobi (version 12.0.1), dADMM and ES-iALM for solving large-scale QROT problems. For our ripALM, we set $\rho=0.99$ based on the numerical observations from the previous section. For ES-iALM, we adopt the same settings for $\rho$, $\sigma_{k}$}, and the termination conditions as those used in ripALM. Moreover, we employ FISTA [3] to solve the ALM-subproblems, with the maximum number of total inner iterations set to 15000.555Note that since ES-iALM does not involve a proximal term, the Ssn method may not be applicable to its subproblems, as the nonsingularity of the generalized Jacobian cannot be guaranteed. Therefore, we employ FISTA as the subsolver in ES-iALM. For Gurobi, we use its default termination conditions and set the corresponding termination tolerances as $10^{-6}$, aligning it with our tolerance. For dADMM, we initialize the penalty parameter as $\sigma_{0}=0.01\|C\|_{F}^{-1}$, and dynamically adjust it based on the primal-dual residuals, following the approach described in [57, Section 5.1], to achieve superior empirical performance. Moreover, we will terminate dADMM when it returns a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})<10^{-6}$, or its number of iterations reaches 10000.
@@
We follow [59, Section 5.1] to generate a random instance. Specifically, we first generate two discrete probability distributions $\big\{(a_{i},\bm{p}_{i})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:i=1,\cdots,m\big\}$ and $\big\{(b_{j},\bm{q}_{j})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:j=1,\cdots,n\big\}$. Here, $\bm{a}:=(a_{1},\cdots,a_{m})^{\top}$ and $\bm{b}:=(b_{1},\cdots,b_{n})^{\top}$ are probabilities/weights, which are generated from the uniform distribution on the open interval $(0,1)$ and further normalized such that $\sum^{m}_{i}a_{i}=\sum^{n}_{j}b_{j}=1$. Moreover, $\{\bm{p}_{i}\}$ and $\{\bm{q}_{j}\}$ are support points whose entries are drawn from a five-component multivariate Gaussian mixture distribution, with a mean vector $(-20,10,0,10,20)^{\top}$ and a variance vector $(5,5,5,5,5)^{\top}$, using randomly generated weights. Then, the cost matrix $C$ is generated by $c_{ij}=\|\bm{p}_{i}-\bm{q}_{j}\|^{2}$ for $1\leq i\leq m$ and $1\leq j\leq n$ and normalized by dividing (element-wise) by its maximal entry.
We then generate a set of random instances with $m=n\in\{1000,\,2000,\dots,10000\}$. For each $m$, we generate 10 instances with different random seeds, and present the average numerical performances of Gurobi, dADMM, ES-iALM, and ripALM in Table 4 and Figure 1, with $\lambda\in\{1,\,0.1\}$. We observe that both ripALM and Gurobi are able to solve the tested problems accurately, in the sense that the residual $\Delta_{\mathrm{res}}^{\rm qrot}$ falls below $10^{-6}$. In contrast, ES-iALM often fails to attain solutions of comparable accuracy, even after reaching its maximum iteration limit and incurring substantially higher computational costs. This is mainly due to the limited efficiency of the first-order subsolver used in ES-iALM. Moreover, dADMM performs even worse: despite consuming a similar amount of computational time as ES-iALM, it yields solutions of considerably lower quality. In addition, we have also observed that Gurobi can be rather time-consuming and memory-consuming for large-scale problems. As an example, for the case where $m=n=10000$, a large-scale QP containing $10^{8}$ nonnegative variables and 20000 equality constraints was solved. One can see that Gurobi is around $2\!\sim\!4$ times slower than our ripALM. In addition, Gurobi may lack robustness, especially for solving large-scale problems. Indeed, as observed from Figure 1, the computational times taken by Gurobi vary a lot among 10 randomly generated instances. In conclusion, our ripALM shows potential for greater efficiency and robustness in solving large-scale QROT problems.
@@
Table 4: Numerical results of Gurobi, dADMM, ES-iALM, and ripALM for solving the QROT problem (4.1). In the table, $\Delta_{\mathrm{res}}^{\rm qrot}$” denotes the terminating $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$; “#” denotes the number of iterations (the total number of Ssn/FISTA iterations is given in the bracket); “time” denotes the computational time.
Gurobi dADMM ES-iALM ripALM
$m=n$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
@@
1000 3.89e-07 32 4.63 3.21e-05 10000 16.11 8.33e-05 16 (15000) 18.87 6.44e-07 18 (67) 1.93
2000 2.71e-07 36 21.88 1.26e-04 10000 102.04 7.87e-05 15 (15000) 107.03 6.49e-07 18 (77) 9.69
3000 3.88e-07 40 57.76 3.10e-04 10000 243.63 8.38e-05 15 (15000) 255.45 6.73e-07 19 (86) 27.50
4000 3.81e-07 41 120.33 5.90e-04 10000 439.32 6.54e-05 14 (15000) 461.86 6.81e-07 19 (93) 56.32
5000 5.93e-07 47 201.37 7.22e-04 10000 689.08 3.25e-05 15 (15000) 729.19 6.18e-07 19 (99) 97.73
6000 5.47e-07 48 308.46 1.12e-03 10000 994.07 3.19e-04 14 (15000) 1047.92 6.50e-07 20 (110) 155.45
7000 4.46e-07 46 409.37 1.40e-03 10000 1354.34 1.67e-04 14 (15000) 1420.33 6.10e-07 20 (116) 225.95
8000 1.37e-06 47 615.47 1.80e-03 10000 1762.95 3.67e-04 14 (15000) 1852.26 6.25e-07 20 (120) 306.10
9000 8.38e-07 50 779.09 2.04e-03 10000 2230.05 1.84e-04 14 (15000) 2343.77 6.82e-07 20 (127) 408.59
10000 7.11e-07 52 1051.46 1.99e-03 10000 2750.64 1.44e-04 14 (15000) 2897.32 7.40e-07 20 (128) 504.35
Figure 1: Comparisons among Gurobi, dADMM, ES-iALM, and ripALM for solving the QROT problem (4.1) with $m=n\in\{1000,\,2000,\cdots,10000\}$. The shaded region indicates the maximum and minimum computational times taken to solve 10 instances for each problem dimension.
5.3 Experiments for BPDN
In this section, we compare our ripALM with Mosek (version 11.0.27), SPGL1 (version 2.1), and dADMM (see Appendix E) for solving the BPDN problem (4.9). For ripALM, we set $\rho=0.99$ as in the previous section. For dADMM, we initialize the penalty parameter as $\sigma=1$, and dynamically update it based on the primal-dual residuals, following the strategy in [57, Section 5.1]. We terminate dADMM when it returns a point $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})<10^{-6}$ or reaches 10000 iterations. For SPGL1, we use its default termination conditions, set the tolerance as $10^{-6}$, (aligning it with our tolerance), and set the maximum number of iterations as 10000. For Mosek, we use its default termination conditions, but adopt two tolerance levels: a tolerance of $10^{-6}$ (denoted as Mosek (1e-6)) to match our tolerance and a tighter tolerance of $10^{-10}$ (denoted as Mosek (1e-10)) to generate a high-precision solution to use as a benchmark.
We construct the tested BPDN problems using both synthetic datasets and UCI datasets [^6] Available at: https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/regression.html. The synthetic datasets are generated as follows. Given a triple $(m,n,s)$, we first generate a dictionary matrix $D\in\mathbb{R}^{m\times n}$ with i.i.d. standard Gaussian entries. We then uniformly select an index set $\mathcal{S}\subset\{1,\cdots,n\}$ with cardinality $|\mathcal{S}|=s$ and construct an $s$-sparse signal $\bar{\bm{s}}\in\mathbb{R}^{n}$, whose entries on $\mathcal{S}$ are drawn independently from the standard Gaussian distribution and are zero elsewhere. The measurement vector is then formed as $\bm{b}=D\bar{\bm{s}}+\delta\bm{\zeta}$, where $\bm{\zeta}\in\mathbb{R}^{m}$ is a noise vector with i.i.d. standard Gaussian entries and $\delta>0$ is the noise level. For UCI datasets, we choose housing7 and mpg7 to obtain $(D,\bm{b})$, where $D$ has dimensions $506\times 77520$ for housing7 and $392\times 3432$ for mpg7. Note that housing7 and mpg7 are expanded versions of their datasets, with the suffix ‘7’ indicating that an order $7$ polynomial is used to generate the basis functions; see [22, Section 4.1] for more details. Finally, to evaluate the solution quality, we compute the relative feasibility residual and normalized objective gap, defined respectively as:
$\texttt{feas}(\bm{s}):=\frac{\max\left\{\|D\bm{s}-\bm{b}\|_{2}-\widehat{\kappa},\,0\right\}}{1+\|\bm{b}\|_{2}},\quad\text{ and }\quad\texttt{nobj}(\bm{s}):=\frac{\big|\|\bm{s}\|_{1}-\|\bm{s}^{*}\|_{1}\big|}{1+\|\bm{s}^{*}\|_{1}},$
where $\bm{s}^{*}$ denote the solution returned by Mosek (1e-10).
Table 5 reports the average computational results for (m,n,s)=(0.1i,i, 0.02i)(m,n,s)=(0.1i,\,i,\,0.02i) with $i\in\{10000,20000,30000\}$ over $10$ randomly generated BPDN instances. In these experiments, the parameter $\widehat{\kappa}$ in (4.9) is set to $\widehat{\kappa}=\delta\left\|\bm{\zeta}\right\|_{2}$ with $\delta\in\{0.01,\,0.1\}$. The results demonstrate the superior performance of ripALM: it consistently achieves the smallest normalized objective gap, while requiring the least computational time and maintaining favorable feasibility residuals. Table 6 reports the results on two UCI data sets, where we choose $\widehat{\kappa}=\delta\|\bm{b}\|$ with $\delta\in\{0.01,\,0.1\}$. These results further show the encouraging performance of our ripALM on the real-world datasets.
Table 5: Numerical results of Mosek, SPGL1, dADMM, and ripALM for solving the BPDN problem (4.9) with $(m,n,s)=(0.1i,i,0.02i)$ and $\widehat{\kappa}=\delta\left\|\bm{\zeta}\right\|_{2}$. In the table, “nobj” denotes the normalized objective gap; “feas” denotes the relative feasibility residual; “#” denotes the number of iterations (the total number of Ssn iterations is given in the bracket); “time” denotes the computational time.
$\delta=0.01$ $\delta=0.1$
$i$ method nobj feas # time nobj feas # time
10000 Mosek (1e-10) 0 7.94e-12 26 6.80 0 1.04e-11 24 6.31
Mosek (1e-6) 1.09e-05 3.42e-07 19 5.52 1.00e-05 3.12e-07 18 5.21
SPGL1 4.72e-05 0 4520 28.32 1.35e-05 0 703 6.15
dADMM 7.49e-06 4.49e-12 10000 71.65 1.95e-06 8.98e-13 4081 31.20
ripALM 1.35e-08 1.31e-09 9 (58) 2.26 4.35e-09 2.89e-09 9 (46) 1.87
20000 Mosek (1e-10) 0 7.50e-12 32 36.20 0 8.51e-12 28 31.61
Mosek (1e-6) 1.12e-05 3.42e-07 21 25.79 1.05e-05 3.20e-07 20 25.02
SPGL1 4.83e-05 0 7991 219.83 3.71e-05 0 997 34.68
dADMM 8.75e-06 1.83e-12 10000 286.40 1.93e-06 5.38e-13 5391 179.74
ripALM 2.83e-09 7.64e-10 9 (81) 12.63 1.06e-08 3.47e-10 10 (62) 10.17
30000 Mosek (1e-10) 0 9.16e-12 30 83.94 0 7.44e-12 32 87.60
Mosek (1e-6) 8.35e-06 2.51e-07 22 67.40 1.31e-05 3.95e-07 21 64.39
SPGL1 5.28e-05 0 8156 479.54 1.69e-05 0 1038 80.32
dADMM 9.36e-06 2.06e-12 10000 601.85 1.88e-06 2.64e-13 6121 453.56
ripALM 5.28e-09 7.82e-10 8 (81) 32.52 4.49e-08 3.88e-10 9 (71) 28.50
Table 6: Same as Table 5 but for the UCI datasets with $\widehat{\kappa}=\delta\|\bm{b}\|$.
$\delta=0.01$ $\delta=0.1$
datasets method nobj feas # time nobj feas # time
housing7 Mosek (1e-10) 0 2.37e-11 36 44.94 0 2.40e-12 27 35.05
Mosek (1e-6) 1.83e-05 4.60e-08 22 25.42 4.31e-05 1.44e-08 19 23.89
SPGL1 1.76e-01 1.42e-02 10000 254.94 9.21e-03 0 1539 51.72
dADMM 4.98e-06 0 10000 237.17 1.74e-06 0 3201 76.53
ripALM 6.58e-06 0 22 (74) 2.28 2.08e-09 0 18 (45) 1.15
mpg7 Mosek (1e-10) 0 1.49e-09 26 1.03 0 1.52e-13 16 0.65
Mosek (1e-6) 1.45e-07 2.98e-08 17 0.62 8.92e-06 1.82e-08 13 0.53
SPGL1 8.82e-01 3.85e-02 10000 3.00 5.35e-03 0 443 0.22
dADMM 1.48e-04 0 10000 3.05 2.33e-06 0 2001 0.60
ripALM 1.03e-06 1.73e-08 30 (126) 0.49 2.06e-07 1.58e-10 18 (23) 0.02